# Corrections to Kerr-Newman from Noncommutative Einstein-Maxwell equation

### v1

@@ Abstract
In this letter we introduce the noncommutative geometry into the standard Einstein-Hilbert-Maxwell action via the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist and solve the equation of motion pertubatively in the expansion of the noncommutative parameter $a$. The equation of motion, the NC Einstein-Maxwell equation, turns out to be effectively a problem in nonlinear electrodynamics where the energy-momentum tensor $T_{\mu\nu}$ obtains correction terms with three Faraday tensors $F_{\mu\nu}$. A solution with nonzero $a^{1}$ terms turns out to be the Kerr-Newman black hole modified with nonzero $g_{t\theta},g_{r\phi},g_{tr}$ and $g_{\phi\theta}$ components proportional to $a$, while the electromagnetic potential is the Seiberg-Witten expanded Kerr-Newman potential which introduces a nonzero $A_{\theta}$ term proportional to $a$.
@@ I Introduction
Noncommutative (NC) geometry is a candidate framework for probing the effects of quantum spacetime in physics. It has been studied in many papers, concerning various topics like black hole entropy [1, 2, 3], gravitational waves [4, 5], gauge theory [6, 7, 8], vacuum energy [9, 10], etc.. In this letter, we will use the Hopf algebra formalism, as outlined in [11, 2, 4] and many other references, to deform the Einstein-Hilbert-Maxwell (EHM) action and solve its equations of motion under the assumption that the solution fields have two Killing vectors, $\partial_{t}$ and $\partial_{\phi}$. We conclude that there exist infinitely many deformations of the EHM action, but all of them share a static solution which we found. In short, the (commutative) geometry of a (commutative) manifold $\mathcal{M}$ is encoded in its Hopf Algebra of vector fields $(\mathcal{U}(\Xi),m,\Delta,\epsilon,S)$, where
@@
and this deformed Hopf algebra is said to encode the noncommutative geometry of some noncommutative manifold $\mathcal{M}_{\star}$. Furthermore, it is also well known that the algebra $(A,\star)$ modules, i.e., the algebras on which the Hopf algebra acts  [^2] All of the relevant geometric modules on which the vector field Hopf algebra naturally acts can be seen as algebras - e.g., the pointwise algebra of functions on the manifold $C^{\infty}(\mathcal{M},\mathbb{C})$, the tensor algebra $\mathcal{T}$, etc. are all algebras., of the deformed Hopf algebra can be obtained from the underlying undeformed Hopf algebra’s algebra modules $(A,\cdot)$ simply by setting  [^3] Where $a_{1}\cdot a_{2}$ can be understood as a tensor product map $\cdot:A\otimes A\rightarrow A$ with $\cdot(a_{1}\otimes a_{2})=a_{1}\cdot a_{2}$.
@@
holds for all $v\in\mathcal{U}(\Xi)$ and all $a_{1},a_{2}\in A$. We will use the the construction (2) and its consequence (3) to construct an action functional that is covariant under the deformed Hopf algebra obtained by deforming the $\mathcal{U}(\Xi)$ Hopf algebra with the $\partial_{t}\wedge\partial_{\phi}$ Drinfeld twist
$\mathcal{F}=e^{-i\frac{a}{2}\left(\partial_{t}\otimes\partial_{\phi}-\partial_{\phi}\otimes\partial_{t}\right)}=e^{-i\frac{a}{2}\partial_{t}\wedge\partial_{\phi}}\;.$ (4)
@@ II Noncommutative Einstein-Hilbert-Maxwell action
In this subsection we will obtain the noncommutative Einstein-Hilbert-Maxwell action. Let us consider the deformed Hopf algebra $(\mathcal{U}(\Xi),m,\Delta_{\mathcal{F}},\epsilon,S_{\mathcal{F}})$ of vector fields over spacetime $\mathcal{M}$ and its deformed algebra module of spacetime functions $C^{\infty}(\mathcal{M},\mathbb{C})$, defined using the Drinfeld twist (4) as above. An example action in this framework is, e.g., the NC charged scalar field and electromagnetism action on the fixed Reissner–Nordström (RN) background defined in [6]:
@@
$\begin{split}\left(f\star g\right)(x)=&\cdot\left[\mathcal{F}^{-1}\left(f\otimes g\right)\right]\\ =&\cdot\left[e^{i\frac{a}{2}\partial_{t}\wedge\partial_{\phi}}\left(f\otimes g\right)\right]\\ =&f(x)\cdot g(x)+i\frac{a}{2}\left[\partial_{t}f(x)\partial_{\phi}g(x)\right.\\ &\left.-\partial_{\phi}f(x)\partial_{t}g(x)\right]+\mathcal{O}(a^{2})\;,\end{split}$ (8)
which is a direct application of (2) to the Drinfeld twist (4). It is important to comment, because it will be key to including the Einstein-Hilbert Lagrangian into the action (6), that in (6) the $\star$ product between metric tensors was replaced with a standard commutative product. That is because $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors of the RN background and as such any function of the metric obeys
@@
due to (8). Additionally, since the fields in (6) are not elements of the algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\cdot)$, but of the deformed algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\star)$, the usual commutative $U(1)$ gauge symmetry is not a symmetry of (6). The only type of gauge transformations that the action (6) can be invariant to is the noncommutative $U(1)$ gauge transformation, denoted as $U(1)_{\star}$, whose infinitesimal expansion is given as
@@
where $\mathcal{L}^{\star}_{G}$ and $\mathcal{L}^{\star}_{EM}$ are the NC Lagrangian densities for the gravitational and electromagnetic sector and they need to reproduce the usual Einstein-Hilbert and Maxwell Lagrangian densities in the commutative limit. The problem is that there are infinitely many such densities. For example, both $\mathcal{L}^{\star(1)}_{EM}$ and $\mathcal{L}^{\star(2)}_{EM}$
$\begin{split}\mathcal{L}^{\star(1)}_{EM}=\sqrt{-g}\star g^{\mu\nu}\star g^{\rho\sigma}\star\hat{F}_{\mu\rho}\star\hat{F}_{\nu\sigma}\;\\ \mathcal{L}^{\star(2)}_{EM}=\sqrt{-g}\star g^{\mu\nu}\star\hat{F}_{\mu\rho}\star g^{\rho\sigma}\star\hat{F}_{\nu\sigma}\;\end{split}$ (14)
reproduce the Maxwell Lagrangian in the commutative limit. So far, the author is not aware of any criterion to prefer any one NC Lagrangian density over all other candidates. A similar problem is present in the NC gauge theory of gravity  [^4] Paper with T. Jurić and N. Kumara Athithamoole to appear on ArXiV soon where infinitely many noncommutative generalizations of geometric quantities, e.g, the NC metric, curvature scalars etc., have the correct commutative limits, but it is impossible to prefer any one choice over the others. But, it is still possible to find a solution to the equations of motion of the action (13). Namely, since all of the NC corrections to the action (13) come from the star product (8) and from the SW map (11), a solution $(g,A)$ of the expanded action whose fields have Killing vectors $\partial_{t}$ and $\partial_{\phi}$, has to be the same as the solution of the action
@@
$=\int d^{4}x\sqrt{-g}\biggl{[}R-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{q}{2}\theta^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}F^{\mu\nu}\biggr{]}+\mathcal{O}(a^{2})\;.$
with the same kind of symmetry in $g$ and $A$. This is due to the property (9) applied to $A$ and $g$. The action (15), which we refer to as the NC Einstein-Hilbert-Maxwell (NC EHM) action, can be thought of as the restriction of the action functional (13) to the part of the classical configuration space where the fields have Killing vectors $\partial_{t}$ and $\partial_{\phi}$. Additionally, the electromagnetic part of the action (15) expands to (12) (with $F_{\alpha\beta}$ terms vanishing due to the time and axial symmetry), so (15) is a nonlinear electrodynamics theory whose nonlinearity is induced by spacetime noncommutativity/quantumness. In conclusion, in a special symmetrical regime it is possible to actually determine the form of the NC action (13). Now it is just a question whether a solution with $\partial_{t}$ and $\partial_{\phi}$ Killing vectors exists for NC EHM action (15). The answer is that it does and we will find it in the following sections.
@@ III The NC Einstein-Maxwell equation
Having obtained the noncommutative Einstein-Hilbert-Maxwell action, we can vary it and obtain the equations of motion for $g_{\mu\nu}$ and $A_{\mu}$, keeping in mind that any solution that we find has to obey the condition that $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors for the metric and for the potential. The equation of motion for the metric, up to $a^{1}$ order, is the NC Einstein equation
@@
We will approach solving (16) and (18) perturbatively in an expansion in $a$, disregarding $a^{2}$ effects. From the form of the energy-momentum tensor (17) and of the Maxwell equation (18), we can see that the commutative solutions to the Einstein-Maxwell problem do not solve the NC problem  [^5] Unless some accidental cancellation of seemingly unrelated terms in (16) and (18) happens which would keep the equations of motion robust to noncommutative corrections. That is not the case for Kerr-Newmann perturbations and there remain nonzero $a^{1}$ terms which are not solved with a purely commutative solution. in case the electromagnetic potential $A_{\mu}$ has $t$ and $\phi$ components, i.e., the electrovacua which describe spaces with both nonzero magnetic and electric fields are described with modified equations of motion. Having said that, we will search for a solution $(g,A)$ which is a perturbation of the Kerr-Newmann (KN) black hole (without assuming slow rotation)
@@
into the NC Einstein and NC Maxwell equations we find the $a^{1}$ source terms for the $h_{\mu\nu}$ and $B_{\mu}$ as the evaluation of the NC equations (16) and (18) on the commutative Kerr-Newmann solution, i.e., as the part of the equations which the KN electrovacuum (20) does not solve. In the NC Maxwell equation (18), $\lambda=r,\phi,\theta$ equations are not solved by the KN commutative ansatz and there are around 50 terms, given as rational functions, in each nonzero component. For the NC Einstein equation we find that there are only 4 simple distinct source term components arising from plugging in the Kerr-Newmann solution into (16):
$G_{\mu\nu}(g^{\text{KN}})-2T_{\mu\nu}(g^{\text{KN}})=aS_{\mu\nu}=\left[\begin{matrix}0&aS_{tr}&aS_{t\theta}&0\\ aS_{rt}&0&0&aS_{r\phi}\\ aS_{\theta t}&0&0&aS_{\theta\phi}\\ 0&aS_{\phi r}&aS_{\phi\theta}&0\end{matrix}\right]$ (22)
@@
$\begin{gathered}S_{tr}=S_{rt}=-\frac{4JQ^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\sin{\left(2\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{r\phi}=S_{\phi r}=\frac{8J^{2}Q^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\left(-2\Sigma{\left(r,\theta\right)}\cos{\left(\theta\right)}+\sin{\left(\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}\right)\sin^{2}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{t\theta}=S_{\theta t}=\frac{4JQ^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{\phi\theta}=S_{\theta\phi}=-\frac{8J^{2}Q^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin^{3}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;.\end{gathered}$ (23)

### v2

# Corrections to Kerr-Newman black hole from noncommutative Einstein-Maxwell equation

@@ Abstract
In this letter we introduce the noncommutative geometry into the standard Einstein-Hilbert-Maxwell action via the $\partial_{t}\wedge\partial_{\varphi}$ Drinfeld twist and solve the equation of motion pertubatively in the expansion of the noncommutative parameter $a$. The equation of motion, the NC Einstein-Maxwell equation, turns out to be effectively a problem in nonlinear electrodynamics where the energy-momentum tensor $T_{\mu\nu}$ obtains correction terms with three Faraday tensors $F_{\mu\nu}$. A solution with nonzero $a^{1}$ terms turns out to be the Kerr-Newman black hole modified with nonzero $g_{t\theta},g_{r\varphi},g_{tr}$ and $g_{\varphi\theta}$ components proportional to $a$, while the electromagnetic potential is the Seiberg-Witten expanded Kerr-Newman potential which introduces a nonzero $A_{\theta}$ term proportional to $a$.
@@ I Introduction
Noncommutative (NC) geometry is a candidate framework for probing the effects of quantum spacetime in physics. It has been studied in many papers, concerning various topics like black hole entropy [1, 2, 3], gravitational waves [4, 5], gauge theory [6, 7, 8], vacuum energy [9, 10], etc.. In this letter, we will use the Hopf algebra formalism, as outlined in [11, 2, 4] and many other references, to deform the Einstein-Hilbert-Maxwell (EHM) action and solve its equations of motion under the assumption that the solution fields have two Killing vectors, $\partial_{t}$ and $\partial_{\varphi}$. We conclude that, in the $\partial_{t}\wedge\partial_{\varphi}$ twist, there exist infinitely many deformations of the EHM action, but all of them share a static solution which we found.
This investigation is related to the current topics in noncommutative geometry in the sense that it uses the same Drinfeld twist approach to noncommutative geometry as the recent papers [1, 2, 3, 4, 5, 6, 7, 11] (and many more) as well as the well established Seiberg-Witten map approach to noncommutative gauge theory [12], but offering an inspection of new, previously unconsidered, phenomenological qualities. Namely, this is the first theoretical observation of a corrected black hole metric arising from the effects of noncommutative geometry as the solution of the equation of motion for some NC action written in NC geometry. Additionally, this is the first investigation of the Kerr-Newman black hole in this framework. The corrections obtained are significant both as a proof of concept, demonstrating that physics in noncommutative geometry can ambitiously pursue nontrivial solutions to the noncommutative Einstein equation, and from a phenomenological perspective, since the corrected metric governs all gravitational properties within the solved system.
In short, the (commutative) geometry of a (commutative) manifold $\mathcal{M}$ is encoded in its Hopf Algebra of vector fields $(\mathcal{U}(\Xi),m,\Delta,\epsilon,S)$, where
@@
and this deformed Hopf algebra is said to encode the noncommutative geometry of some noncommutative manifold $\mathcal{M}_{\star}$. Furthermore, the algebra modules, i.e., the algebras on which the Hopf algebra acts  [^2] All of the relevant geometric modules on which the vector field Hopf algebra naturally acts can be seen as algebras - e.g., the pointwise algebra of functions on the manifold $C^{\infty}(\mathcal{M},\mathbb{C})$, the tensor algebra $\mathcal{T}$, etc. are all algebras., of the twisted Hopf algebra are well known. Any algebra $(A,\star)$ module of the deformed Hopf algebra can be obtained from the underlying undeformed Hopf algebra’s module $(A,\cdot)$ simply by setting  [^3] Where $a_{1}\cdot a_{2}$ can be understood as a tensor product map $\cdot:A\otimes A\rightarrow A$ with $\cdot(a_{1}\otimes a_{2})=a_{1}\cdot a_{2}$.
@@
holds for all $v\in\mathcal{U}(\Xi)$ and all $a_{1},a_{2}\in A$. We will use the construction (2) and its consequence (3) to construct an action functional that is covariant under the deformed Hopf algebra obtained by deforming the $\mathcal{U}(\Xi)$ Hopf algebra with the $\partial_{t}\wedge\partial_{\varphi}$ Drinfeld twist
$\mathcal{F}=e^{-i\frac{a}{2}\left(\partial_{t}\otimes\partial_{\varphi}-\partial_{\varphi}\otimes\partial_{t}\right)}=e^{-i\frac{a}{2}\partial_{t}\wedge\partial_{\varphi}}\;.$ (4)
@@ II Noncommutative Einstein-Hilbert-Maxwell action
In this section we will obtain the noncommutative Einstein-Hilbert-Maxwell action. Let us consider the deformed Hopf algebra $(\mathcal{U}(\Xi),m,\Delta_{\mathcal{F}},\epsilon,S_{\mathcal{F}})$ of vector fields over spacetime $\mathcal{M}$ and its deformed algebra module of spacetime functions $C^{\infty}(\mathcal{M},\mathbb{C})$, defined using the Drinfeld twist (4) as above. An example action in this framework is, e.g., the NC charged scalar field and electromagnetism action on the fixed Reissner–Nordström (RN) background defined in [6]:
@@
$\begin{split}\left(f\star g\right)(x)=&\cdot\left[\mathcal{F}^{-1}\left(f\otimes g\right)\right]\\ =&\cdot\left[e^{i\frac{a}{2}\partial_{t}\wedge\partial_{\varphi}}\left(f\otimes g\right)\right]\\ =&f(x)\cdot g(x)+i\frac{a}{2}\left[\partial_{t}f(x)\partial_{\varphi}g(x)\right.\\ &\left.-\partial_{\varphi}f(x)\partial_{t}g(x)\right]+\mathcal{O}(a^{2})\;,\end{split}$ (8)
which is a direct application of (2) to the Drinfeld twist (4). It is important to comment, because it will be key to including the Einstein-Hilbert Lagrangian into the action (6), that in (6) the $\star$ product between metric tensors was replaced with a standard commutative product. That is because $\partial_{t}$ and $\partial_{\varphi}$ are Killing vectors of the RN background and as such any function of the metric obeys
@@
due to (8). Additionally, since the fields in (6) are not elements of the algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\cdot)$, but of the deformed algebra $(C^{\infty}(\mathcal{M},\mathbb{C}),\star)$, the usual commutative $U(1)$ gauge symmetry is not a symmetry of (6). The only type of gauge transformation that the action (6) can be invariant to is the noncommutative $U(1)$ gauge transformation, denoted as $U(1)_{\star}$, whose infinitesimal expansion is given as
@@
where $\mathcal{L}^{\star}_{G}$ and $\mathcal{L}^{\star}_{EM}$ are the NC Lagrangian densities for the gravitational and electromagnetic sector and they need to reproduce the usual Einstein-Hilbert and Maxwell Lagrangian densities in the commutative limit.
In this paper, we consider only the minimally NC deformed Lagrangians which only replace pointwise commutative $\cdot$ products with $\star$ products and gauge fields with Seiberg-Witten expanded NC gauge fields in the Einstein-Hilbert-Maxwell Lagrangian density. The problem is that there are infinitely many such densities. For example, $\mathcal{L}^{\star(1)}_{EM}$, $\mathcal{L}^{\star(2)}_{EM}$ and $\mathcal{L}^{\star(3)}_{EM}$
$\begin{split}\mathcal{L}^{\star(1)}_{EM}&=\sqrt{-g}\star g^{\mu\nu}\star g^{\rho\sigma}\star\hat{F}_{\mu\rho}\star\hat{F}_{\nu\sigma}\;\\ \mathcal{L}^{\star(2)}_{EM}&=\sqrt{-g}\star g^{\mu\nu}\star\hat{F}_{\mu\rho}\star g^{\rho\sigma}\star\hat{F}_{\nu\sigma}\;\\ \mathcal{L}^{\star(3)}_{EM}&=-2\mathcal{L}^{\star(1)}_{EM}+3\mathcal{L}^{\star(2)}_{EM}\end{split}$ (14)
all reproduce the Maxwell Lagrangian in the commutative limit, but they all give rise to mutually inequivalent action functionals. When talking about NC deformations of a theory, the author means choosing one such NC Einstein-Hilbert-Maxwell Lagrangian arising from the twist (4) and the NC geometry and gauge theory frameworks outlined in Sections I and II  [^4] One can imagine also adding some fundamentally new terms to the Einstein-Hilbert-Maxwell action proportional to the NC parameter $a$. For example, $aR^{2}$ or almost anything else. Such actions would also revert back to the standard action in the commutative limit. But, in this paper we only consider the deformations which arise from submerging the Einstein-Hilbert-Maxwell action into the NC geometry and NC gauge theory framework without adding any new terms.. From now on, we will suppose that (13) are minimal NC deformations in the sense that was just described. So far, the author is not aware of any criterion to prefer any one NC Lagrangian density over all other candidates. A similar problem is present in the NC gauge theory of gravity [17] where infinitely many noncommutative generalizations of geometric quantities, e.g, the NC metric, curvature scalars etc., have the correct commutative limits, but it is impossible to prefer any one choice over the others. But, it is still possible to find a solution to the equations of motion of the action (13). Namely, since all of the NC corrections to the action (13) come from the star product (8) and from the SW map (11), a solution $(g,A)$ of the expanded action whose fields have Killing vectors $\partial_{t}$ and $\partial_{\varphi}$, has to be the same as the solution of the action
@@
$=\int d^{4}x\sqrt{-g}\biggl{[}R-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{q}{2}\theta^{\alpha\beta}F_{\mu\alpha}F_{\nu\beta}F^{\mu\nu}\biggr{]}+\mathcal{O}(a^{2})$
with the same kind of symmetry in $g$ and $A$. This is due to the property (9) applied to $A$ and $g$. The action (15), which we refer to as the NC Einstein-Hilbert-Maxwell (NC EHM) action, can be thought of as the restriction of the action functional (13) to the part of the classical configuration space where the fields have Killing vectors $\partial_{t}$ and $\partial_{\varphi}$. Additionally, the electromagnetic part of the action (15) expands to (12) (with $F_{\alpha\beta}$ terms vanishing due to the time and axial symmetry), so (15) is a nonlinear electrodynamics theory whose nonlinearity is induced by spacetime noncommutativity/quantumness. In conclusion, in a special symmetrical regime it is possible to actually determine the form of the NC action (13) without specifying the ordering of $\star$ products. Now it is just a question whether a solution with $\partial_{t}$ and $\partial_{\varphi}$ Killing vectors exists for NC EHM action (15). The answer is that it does and we will find it in the following sections.
@@ III The NC Einstein-Maxwell equation
Having obtained the noncommutative Einstein-Hilbert-Maxwell action, we can vary it and obtain the equations of motion for $g_{\mu\nu}$ and $A_{\mu}$, keeping in mind that any solution that we find has to obey the condition that $\partial_{t}$ and $\partial_{\varphi}$ are Killing vectors for the metric and for the potential. The equation of motion for the metric, up to $a^{1}$ order, is the NC Einstein equation
@@
We will approach solving (16) and (18) perturbatively in an expansion in $a$, disregarding $a^{2}$ effects. From the form of the energy-momentum tensor (17) and of the Maxwell equation (18), we can see that the commutative solutions to the Einstein-Maxwell problem do not solve the NC problem  [^5] Unless some accidental cancellation of seemingly unrelated terms in (16) and (18) happens which would keep the equations of motion robust to noncommutative corrections. That is not the case for Kerr-Newmann perturbations and there remain nonzero $a^{1}$ terms which are not solved with the purely commutative solution. in case the electromagnetic potential $A_{\mu}$ has $t$ and $\varphi$ components, i.e., the electrovacua which describe spaces with both nonzero magnetic and electric fields are described with modified equations of motion. Having said that, we will search for a solution $(g,A)$ which is a perturbation of the Kerr-Newmann (KN) black hole (without assuming slow rotation)
@@
into the NC Einstein and NC Maxwell equations we find the $a^{1}$ source terms for the $h_{\mu\nu}$ and $B_{\mu}$ as the evaluation of the NC equations (16) and (18) on the commutative Kerr-Newmann solution, i.e., as the part of the equations which the KN electrovacuum (20) does not solve. In the NC Maxwell equation (18), $\lambda=r,\varphi,\theta$ equations are not solved by the KN commutative ansatz and there are around 50 terms, given as rational functions, in each nonzero component. For the NC Einstein equation we find that there are only 4 simple distinct source term components arising from plugging in the Kerr-Newmann solution into (16):
$G_{\mu\nu}(g^{\text{KN}})-2T_{\mu\nu}(g^{\text{KN}})=aS_{\mu\nu}=\left[\begin{matrix}0&aS_{tr}&aS_{t\theta}&0\\ aS_{rt}&0&0&aS_{r\varphi}\\ aS_{\theta t}&0&0&aS_{\theta\varphi}\\ 0&aS_{\varphi r}&aS_{\varphi\theta}&0\end{matrix}\right]$ (22)
@@
$\begin{gathered}S_{tr}=S_{rt}=-\frac{4JQ^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\sin{\left(2\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{r\varphi}=S_{\varphi r}=\frac{8J^{2}Q^{3}r^{2}\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)\left(-2\Sigma{\left(r,\theta\right)}\cos{\left(\theta\right)}+\sin{\left(\theta\right)}\frac{\partial}{\partial\theta}\Sigma{\left(r,\theta\right)}\right)\sin^{2}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{t\theta}=S_{\theta t}=\frac{4JQ^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin{\left(2\theta\right)}}{M\Sigma^{6}{\left(r,\theta\right)}}\;,\\ S_{\varphi\theta}=S_{\theta\varphi}=-\frac{8J^{2}Q^{3}r\left(r\frac{\partial}{\partial r}\Sigma{\left(r,\theta\right)}-\Sigma{\left(r,\theta\right)}\right)^{2}\Delta{\left(r\right)}\sin^{3}{\left(\theta\right)}\cos{\left(\theta\right)}}{M^{2}\Sigma^{6}{\left(r,\theta\right)}}\;.\end{gathered}$ (23)